What mass of KCl would be recovered if 55 mL of 0.20 M KCl were “evaporated to dryness”?
[Hint: this is the same as asking how many grams of KCl are in 55 mL of 0.20 M KCl.]
1
Expert's answer
2020-10-01T07:51:45-0400

According to the equation:

c = n / V

where c - molar concentration, n - number of moles, V - volume.

From here:

n = c × V

n (KCl) = c × V = 0.20 M × 55 mL = 0.20 M × 0.055 L = 0.011 mol

The mass of the KCl can be calculated by the following formula:

m = n × Mr

where m - mass, n - number of moles, Mr - molecular weight.

From here:

m (KCl) = n × Mr = 0.011 mol × 74.55 g/mol = 0.82 g


Answer: 0.82g